Find the slope of the line containing the given pair of points. If the line is undefined, state so.
(5, 3) and (8, -4)

Answer by mathstutor458(57)

slope= (y2-y1)/(x2-x1)
(x1,y1)=(5,3)
(x2,y2)=(8,-4)
slope=(-4-3)/(8-5) = -7/3 --->answer